OpenStudy (henrietepurina):

Algebra 1 question :D Medal + Fan!

OpenStudy (henrietepurina):

A quadratic equation is shown below: 9x^2 - 36x + 36 = 0 Describe the solution(s) to the equation by just determining the radicand. Show your work.

OpenStudy (anonymous):

divide by 9 first

OpenStudy (henrietepurina):

ok gimme a sec :D

OpenStudy (henrietepurina):

x^2−4x+4

OpenStudy (henrietepurina):

(refresh page)

OpenStudy (henrietepurina):

would that be correct @satellite73 ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[x^2-4x+4=0\] is the start

OpenStudy (anonymous):

then it is a perfect square, namely \((x-2)^2=0\)

OpenStudy (henrietepurina):

ok then what should I do next...

OpenStudy (anonymous):

say if \(x-2=0\) then \(x=2\) and be done with it

OpenStudy (henrietepurina):

ok then, and was that the solution by just determining the radicand? @satellite73

OpenStudy (henrietepurina):

also @satellite73 , why did you go from (x−2)^2=0 to x−2=0?

OpenStudy (anonymous):

because if the square of a number is zero, the number itself must be zero

OpenStudy (henrietepurina):

ah ok, so overall, you determined the solution by just determining the radicand? @satellite73

OpenStudy (anonymous):

i am not sure of this terminology to me the radicand means the number inside the radical

OpenStudy (anonymous):

so i you had \((x-2)^2=0\) i guess you could say \(x-2=\sqrt 0\) and the radicand would be 0

OpenStudy (henrietepurina):

same here, but if that meant determining the number inside the radical, would the answer than be x−2=0√ and the radicand would be 0?

OpenStudy (anonymous):

yes i guess, but that makes \(x-2=0\) you want \(x\) which is \(2\)

OpenStudy (henrietepurina):

ok, mind helping me with one more?

OpenStudy (henrietepurina):

thanks btw :D

OpenStudy (anonymous):

sure no problem

OpenStudy (henrietepurina):

A sandbag was thrown downward from a building. The function f(t) = -16t^2 - 32t + 128 shows the height f(t), in feet, of the sandbag after t seconds. Question: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function.

OpenStudy (anonymous):

we can factor this for sure

OpenStudy (henrietepurina):

ok then...

OpenStudy (anonymous):

first take out the common factor of \(16\)

OpenStudy (anonymous):

or better yet, \\(-16\)

OpenStudy (henrietepurina):

yeah yeah... gimme a sec

OpenStudy (henrietepurina):

wait but how?

OpenStudy (anonymous):

how to factor out the \(-16\)?

OpenStudy (henrietepurina):

yeah

OpenStudy (anonymous):

\[-16t^2-32t+128=-16(something)\]

OpenStudy (anonymous):

it it clear what you get or no? if not i will show you

OpenStudy (henrietepurina):

no no, I can do this :D

OpenStudy (anonymous):

ok then it will be easy to factor

OpenStudy (anonymous):

whoa hold the phone tooooo much work it says factor, not solve the quadratic

OpenStudy (henrietepurina):

ok sorry about that

OpenStudy (anonymous):

\[-16t^2-32t+128=-16(t^2+2t-8)\]is a start

OpenStudy (henrietepurina):

−16t^2−30t+136=−16(t^2)

OpenStudy (henrietepurina):

−15t^2−30t+136=−16

OpenStudy (anonymous):

oh my

OpenStudy (henrietepurina):

what... did I go deep down wrong?

OpenStudy (anonymous):

i will try to say it again \[-16t^2-32t+128\] each term has a common factor that common factor is \(16\) or \(-16\)

OpenStudy (henrietepurina):

OH YES THEY DO I JUST NOTICED :D

OpenStudy (anonymous):

so job number one is to "factor it out" i..e factor out the \(-16\) from each term

OpenStudy (henrietepurina):

(t^2)-2t+8

OpenStudy (henrietepurina):

no wait, sorry (t^2)-2t-8

OpenStudy (anonymous):

actually \[-16(t^2+2t-8)\]

OpenStudy (henrietepurina):

oh right... what next?

OpenStudy (anonymous):

then leave the \(-16\) hanging out front, and see if you can factor \[t^2+2t-8\] which should not be too hard

OpenStudy (henrietepurina):

(x−2)(x+4)

OpenStudy (anonymous):

damn question marks should be \((t-2)(t+4)\) if that is what you have then yes

OpenStudy (henrietepurina):

ikr.... oh sorry, I accidentally replaced t with x :P

OpenStudy (anonymous):

makes no real difference

OpenStudy (henrietepurina):

ok then, so, did we just interpret the meaning of x-intercept in this equation?

OpenStudy (anonymous):

it is is when the height is zero there are two intercepts \((t=2,t=-4\) but since we don't really go back in time, \(t=-4\) is not meaningful in this question, and what this means is it took two seconds to hit the ground

OpenStudy (henrietepurina):

ok then, thank you, and have a good rest of your day :D

OpenStudy (anonymous):

you too, and congratulations on getting in to oxbridge !

OpenStudy (henrietepurina):

how did you know?

OpenStudy (anonymous):

psychic

OpenStudy (henrietepurina):

haha... well thanks again, bye :D (it was facebook wasn't it?)

OpenStudy (anonymous):

yes, it was

OpenStudy (anonymous):

and your welcome any time

OpenStudy (henrietepurina):

:D